you purchase a home for 125.000 using function notation write a funvtion that represent the value of your home t years after you purchased it given the following a) the valueof your home increases at a constant rate of 1150 per year v(t)= Be sureto use the variable t in your answer. dont use any commas (b) in how many years will be the home be worth 133,o50 the home will be worth 133, 050 in years.(c) write a sentence describing what the ordered pair(15,142,250) means in the context of the problem be sure to use the correct units in your answer. the meaning of (15,142,250) is
Answer by Theo(13342)

the home is purchased for 125000
the rate of increase in the value of your home is 1150 per year.
your formula is v(t) = 125000 + 1150 * t
t is the number of years you owned the home.

you want to know how many years before your home reaches 133050 in value.
this means that v(t) = 133050.
your equation becomes 133050 = 125000 + 1150 * t
subtract 125000 from both sides of the equation to get:
8050 = 1150 * t
divide both sides of the equation by 1150 to get:
8050 / 1150 = t
solve for t to get t = 7
your home will be equal to 133050 in value in 7 years.

the ordered pair (15,142250) on a graph in (x,y) format means that the value of the home will be 142250 in 15 years.
x represents the number of y4ears on the graph.
y represents the value of the home on the graph.

your graph would look like this.



the graphing software at https://www.desmos.com/calculator